OpenStudy (anonymous):

you invest 6000 dollars into two accounts paying 6% and 9% annual interest ,respectively At the end of the year the accounts earn the same interest. HOW MUCH WAS INVESTED AT EACH RATE. PLEEEEEEEEASE HELP

OpenStudy (anonymous):

\[.09\times x=.06\times (6000-x)\] where x is amount invested at 9%

OpenStudy (anonymous):

i get \[x=\$2400\]

OpenStudy (anonymous):

I'd just use ratios, 9 = 1.5 * 6, so put the amounts in the opposite ratio. If that makes sense.